OpenStudy (kmullis6):

Show me how to work this out? I'm so confused. /: Solve for x: -3(x + 3) = -3(x + 1) - 5. a) 6 b) -6 c) All real numbers d) No solution

OpenStudy (anonymous):

-3 (x + 3) = -3 (x + 1) -5 X + 3 = (x + 1) -5 (divide both sides by -3) X + 8 [is not equal to] x + 1 (add both sides by 5) Therefore, d) No solution

OpenStudy (texaschic101):

-3(x + 3) = -3(x + 1) - 5 --- distribute through the parenthesis -3x - 9 = -3x - 3 - 5 -3x - 9 = -3x - 8 -- add 3x to both sides -9 = -8 (incorrect) NO SOLUTION